1, Oil with a viscosity coefficient of 15 poise is filled between parallel plates spaced 5 mm apart. When the lower plate is fixed and the upper plate moves at a speed of 5 m/s. what is the shear stress (Pa) generated on the plate surface?

2, When 1200N of lead is suspended from the bottom of cube (wood) of 1 m on a side and placed in water, how many cm is the height of the piece of wood that rises above the water ? (the specific gravity of wood block is 0.4, the specific gravity of lead is 12, and the bottom of the wood block is kept horizontal.



1
Expert's answer
2021-10-30T11:20:20-0400

1)

Space between plates dy=5mm=0.005m

Viscosity=15 poise=1.5Ns/m2

Velocity du=5m/s

We know that shear stress,"\tau=\mu\frac{du}{dy}"

"=1.5*\frac{5}{0.005}"

"=1500N\/m^2"

"=1500Pa"

2)

Wlead=Wwood

"\rho_lA_lhg=\rho_wA_wyg"

"1200=0.4*1000*1^2*y*9.8"

"y=\frac{1200}{0.4*1000*9.8}=\frac{1200}{409.8}"

"=2.928m"

=292.8cm